### v1

# Demonstration of sequential processors with quantum advantage and analysis of classical performance limits

Abstract
In this paper, we theoretically and experimentally analyze sequential processors with limited communication between parts. We compare the expressivity of sequential quantum and classical processors under the same constraints. They consist of three or four modules each of which processes local data. The modules of the quantum processor are linked through one qubit or one qutrit communication, while those of the classical processor communicate through one bit or one trit. For the classical processor, we prove bounds on its performance in terms of inequalities on correlations of the output with a target function. We show that the quantum processor violates these inequalities. We show this violation experimentally on a silicon photonics setup. We describe how to find the classical bound on correlations with arbitrary target function by reducing the problem to minimization of an Ising-type spin-glass Hamiltonian. Our theory is applicable in general problems, such as the low-rank binary matrix approximation.
preprint: APS/123-QED
I Background and overview
Studying and comparing simple structure classical and quantum processors is motivated by the urge to understand the fundamental differences between them. This interest is driven by progress in quantum technology in general and quantum computing in particular. In addition, quick progress in such applications as machine learning raised new questions about the benefits of quantum processing units.
In [1], the authors studied a simple one-qubit processor as an effective classifier. The classical data points $X_{i}$ were encoded in classically tuned parameters of the circuit gates; see Fig. 1. Free parameters of the circuit allowed for training the processor so that the finally measured probability of a chosen state could be used as an indicator of a classification label.
Figure 1: Single qubit and single bit sequential processors of similar structure.
The authors proved the universality of such a classifier, that is, an analogue of the universal approximation theory in a single-layer neural network. This proved that even a single-qubit system can compute complicated functions of parameters of the circuit if it is long enough and is appropriately trained. These ideas were experimentally demonstrated [2, 3, 4, 5].
Universality holds for both a quantum and a classical processors if they are sufficiently extended. In such a situation, a natural question is if there is an advantage of a quantum single-qubit processor over a fairly constructed classical counterpart with fixed finite size. The authors in [1] discussed the possible advantage from entanglement and in [6] from parallelization. In the context of communication complexity [7, 8], the authors formulated a similar problem showing a quantum advantage in sending information in qubit states over the analogous scheme with classical bits. Similar conclusions appeared in the context of temporal entanglement and temporal Bell inequalities [9, 10]. In another paper [11], the author formulated the problem in the context of macro-realism and showed a quantum advantage in speeding-up the computation of certain functions. In [7] an experiment was performed on bulk optics showing that a quantum multiparty communication line with a single qubit of communication (analogous to the sequential processor in Fig. 1) can output a binary target function of local parameters $\{X_{i}\}$ which is not achievable in a single-bit communication classical setup. Also, in [12, 13] the authors showed that certain deterministic outputs from quantum processors cannot be generated by classical circuits in limited space. In this sense, one can speak of a larger expressive power of the quantum processor as in [14, 15].
To show these results, it was necessary to prove rigorous limits on how well some functions could be approximated by the output of classical fixed-structure processors. The functions were shown for which the classical limit was known or established by the authors. However, for arbitrary functions this classical limit is unknown, and a complete understanding of the advantage of the quantum processor over the classical one even in limited space is still a subject of research.
Here, we study some aspects of this problem posing the following questions:
  1. 1.
    Is there an advantage of using a fixed length quantum sequential processor with a single-qubit (or single qutrit) communication between processing modules, with respect to the same structure classical processor processing a single bit (or trit)?
  2. 2.
    Can we quantify it rigorously?
  3. 3.
    Can we demonstrate it experimentally on the silicon photonic platform?
Our original results are summarized as follows.
  • We report an experimental demonstration of the qubit and qutrit sequential processors on silicon photonics chip and demonstrate the achievable advantage of quantum processors in realizing certain functions of the input parameters. The qubit processor demonstration is analogous to the earlier bulk optics experiment [7], however, different technology–silicon photonics, and different degrees of freedom–a single photon in two modes are involved.
  • To face specific conditions of the design of the silicon photonics chip with four stages and qutrit communication (two photons in two modes) we needed to develop the theory showing the smallest Hamming distance between the deterministic part of the ideal quantum processor output and the optimal output of the corresponding classical processor. We established this bound based on the properties of the quantum processor output and optimization in the space of binary functions.
  • We showed that the problem of minimal distance between an arbitrary binary target function and the output of a classical sequential processor with limited communication can be reduced to optimization in the Ising model.
  • We successfully solved the Ising problems from the previous point for the classical processors with some target functions. Our models involved up to 48 spins with multipartite interaction. The models were solved on a simulated annealing machine Fixstars Amplify.
  • We indicate consequences of mapping the problem of the smallest distance between the output of a classical sequential processor and a given target for such problems as binary low-rank matrix completion and low-rank binary tensor approximation.
II Limited communication sequential processor
Consider a sequential processor consisted of several data processing modules, see Fig. 1. In each stage, the module processes the information obtained from the earlier section together with the input provided only locally at this stage $X_{i}$ and unknown to other parts. We assume that the information transferred between the modules is strongly restricted. The last module should output a value that depends on all bits of data $\{X_{i}\}$ inputted at all stages, $O(\{X_{i}\})$. The design of a sequential logic processor consists of optimizing each module so that the output function approaches the desired target function $T(\{X_{i}\})$. The latter can be interpreted as a truth table for a logical function with Boolean variables $\{X_{i}\}$.
In Fig. 1 $a)$ we show a three-module quantum sequential processor with two classical bits $(X_{i},X_{i+1})$ processed in each module and a single-qubit inter-module communication. Fig. 1 $b)$ shows its classical counterpart with a one-bit communication. In the modules of the quantum processor, we allow for unitary transformations as well as measurements and classical functions. Therefore, the quantum processor from Fig. 1 $a)$ is at least as powerful in any task as the corresponding classical processor. In the classical circuit, in addition to the restriction that intermodular communication is limited, there are no additional restrictions on the local functions performed by the modules.
In our research, we consider a specific realization of the quantum processor described in detail in the following sections.
II.1 Quantum one-qubit processor
Consider a specific three-stage quantum processor with single-qubit communication between the modules and two classical bit inputs $(X_{2i},X_{2i+1})$ at each module. Depending on the input bits, we fix one of four one-qubit gates on the qubit obtained from the earlier part:
$(0,0)$ $\rightarrow$ $\hat{1}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$, (1)
$(0,1)$ $\rightarrow$ $\hat{\sigma}_{z}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$, (2)
$(1,0)$ $\rightarrow$ $\hat{H}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&i\\ i&1\end{bmatrix}$, (3)
$(1,1)$ $\rightarrow$ $\hat{H}^{*}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&-i\\ -i&1\end{bmatrix}$. (4)
The initial qubit state is $|0\rangle$. In the end, the qubit is measured in the computational basis $\{|0\rangle,|1\rangle\}$. For all $2^{6}$ input bit values $X_{1},...,X_{6}$ ordered in lexicographical order, the final measurement of our idealized processor returns either deterministic outputs $|0\rangle$ or $|1\rangle$, or a state which is a superposition of the two basis states. We define the target function as follows:
  • $T(\{X_{i}\})=1$ if the output is deterministically $|0\rangle$
  • $T(\{X_{i}\})=-1$ if the output is deterministically $|1\rangle$
  • $T(\{X_{i}\})=0$ if the output state is a superposition of $|0\rangle$ and $|1\rangle$.
The target function, after reshaping into a matrix, for convenience, can be written explicitly as follows,
$T(X_{1}X_{2},X_{3}...X_{6})=\left[\begin{array}[]{cccccccccccccccccccc}-1&1&0&0&1&-1&0&0&0&0&1&-1&0&0&-1&1\\ 1&-1&0&0&-1&1&0&0&0&0&-1&1&0&0&1&-1\\ 0&0&1&-1&0&0&-1&1&1&-1&0&0&-1&1&0&0\\ 0&0&-1&1&0&0&1&-1&-1&1&0&0&1&-1&0&0\end{array}\right]$ (5)
where the first two bits determine the row and the remaining four bits determine the column. This notation can be associated with the Karnaugh maps [16] known in the literature on logic design. Alternatively, it can be seen as a tensor as we refer to it in Section VI.
We used this target function to calculate the correlation with the realistic experimental output. The experiment and the results are presented in Section III. There, based on the experimental output $O(x_{i})$ of $N$ randomly selected sequences $x_{i}$ of six bits $X_{1},...,X_{6}$ we calculate the correlation function
$C=\frac{1}{N}\sum_{i=1}^{N}O(x_{i})T(x_{i}).$ (6)
In Section IV we prove bounds on the achievable correlation function for optimized classical sequential processor with three stages, and one-bit communication to be compared with the experimental results.
II.2 Quantum one-qutrit processor
We do a similar analysis for a quantum sequential processor with 3 and 4 modules and a communication carried forward by a three-level quantum system (qutrit). Our analysis is tailored to a quantum optics experiment on silicon photonics, so the qutrit we consider is realized by a state of two photons in two modes. The Hilbert space of the system is spanned by three vectors in the photon number notation $\rangle, |02|02\rangle$, and $|11\rangle$, where $|20\rangle$ means that there are two photons in the first mode and zero photons in the second, etc. As input state, we consider $|20\rangle$. As gates, depending on each module input bits $(X_{2i},X_{2i+1})$, we take:
  • For the three-stage processor and for modules 2,3, and 4 of the four-stage processor we applied gates (1)-(4) acting on the photon creation operators of the modes.
  • For module 1 of the four step processor, because the part of the circuit of our chip did not allow to apply the gates as in the other modules, we applied the following gates
    $(0,0)$ $\rightarrow$ $\frac{1}{\sqrt{2}}\begin{bmatrix}1&i\\ -i&-1\end{bmatrix}$, (7)
    $(0,1)$ $\rightarrow$ $\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$, (8)
    $(1,0)$ $\rightarrow$ $\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}$, (9)
    $(1,1)$ $\rightarrow$ $\frac{1}{\sqrt{2}}\begin{bmatrix}1&-i\\ i&-1\end{bmatrix}$. (10)
If the final measurement of an ideal processor deterministically indicates the state $|20\rangle$ we define the entry of the corresponding target function as 1. If it deterministically indicates the state $|02\rangle$, we define the entry 1-1. Otherwise, the target function entry is set to 0. The quantum processor we built does not allow us to achieve state $|11\rangle$ deterministically from the initial state $|20\rangle$ by any unitary transformation that preserves the bosonic commutation relations. Therefore, the target function of our qutrit processor resembles a qubit processor target.
The target of the four part processor after reshuffling as a matrix, for convenience, is explicitly given as
$T(X_{1}...X_{4},X_{5}...X_{8})=\left[\begin{array}[]{cccccccccccccccccccc}0&0&-1&1&0&0&1&-1&-1&1&0&0&1&-1&0&0\\ 0&0&1&-1&0&0&-1&1&1&-1&0&0&-1&1&0&0\\ -1&1&0&0&1&-1&0&0&0&0&1&-1&0&0&-1&1\\ 1&-1&0&0&-1&1&0&0&0&0&-1&1&0&0&1&-1\\ 1&-1&0&0&-1&1&0&0&0&0&-1&1&0&0&1&-1\\ -1&1&0&0&1&-1&0&0&0&0&1&-1&0&0&-1&1\\ 0&0&-1&1&0&0&1&-1&-1&1&0&0&1&-1&0&0\\ 0&0&1&-1&0&0&-1&1&1&-1&0&0&-1&1&0&0\\ 0&0&1&-1&0&0&-1&1&1&-1&0&0&-1&1&0&0\\ 0&0&-1&1&0&0&1&-1&-1&1&0&0&1&-1&0&0\\ 1&-1&0&0&-1&1&0&0&0&0&-1&1&0&0&1&-1\\ -1&1&0&0&1&-1&0&0&0&0&1&-1&0&0&-1&1\\ -1&1&0&0&1&-1&0&0&0&0&1&-1&0&0&-1&1\\ 1&-1&0&0&-1&1&0&0&0&0&-1&1&0&0&1&-1\\ 0&0&1&-1&0&0&-1&1&1&-1&0&0&-1&1&0&0\\ 0&0&-1&1&0&0&1&-1&-1&1&0&0&1&-1&0&0\\ \end{array}\right].$ (11)
The three-stage processor has the target function the same as the first $4\times 16$ block of this matrix.
III Silicon photonics experiment
We implemented both a one-qubit and a one-qutrit processor on a silicon photonic integrated circuit and evaluated their performance using correlation functions [17, 18, 19]. As shown in Fig. 2(a), the experimental setup employed a $\times 4$ universal unitary transformation circuit composed of beam splitters and phase shifters [20, 21]. In this study, we used two spatial modes of the circuit to construct a sequential processor by cascading multiple unitary transformations. Each unitary transformation consisted of a phase shifter and a Mach–Zehnder interferometer (MZI), allowing us to realize the target operations with high precision.
Figure 2: (a) Schematic of the sequential quantum processor implemented on a silicon photonic chip. The unitary gate was implemented using a heater and an MZI enclosed by the red dotted lines. (b) Overview of the one-qubit processor. A heralded single photon was injected into the sequential processor. (c) Overview of the one-qutrit processor. Two photons generated from attenuated laser light at the single-photon level were injected into the sequential processor. In Figs. (b) and (c), the circuit enclosed by the red dotted lines corresponds to a single unitary gate in each case.
For the implementation of the one-qubit processor, heralded single photons were employed, as illustrated in Fig. 2(b). Photon pairs were generated via spontaneous four-wave mixing (SFWM), utilizing the third-order nonlinearity of silicon [22, 23]. Specifically, a continuous-wave (CW) pump light with a wavelength of 1549.32nm and an output power of approximately 10mW was injected into a silicon waveguide, generating signal and idler photons through SFWM. The signal photons had a central wavelength of 1559.0 $\pm$ 0.8nm, while the idler photons had a central wavelength of 1539.8 $\pm$ 0.8nm. The spectral width of these photons is determined by the bandwidth of the off-chip frequency filter (200 GHz). The signal photons were used as heralds (triggers), and the idler photons were directed to the processor circuit [24, 25, 26]. After passing through these filters, both the idler and corresponding signal photons were detected using superconducting nanowire single-photon detectors (SNSPDs). The generation rate of signal single photons obtained after passing through the circuit was approximately 30 counts per second.
In this experiment, the input state was $|1;0\rangle_{ab}$, where a single photon was present in mode $a$. Randomly selected operations from the set $\{\hat{I},\hat{\sigma}_{z},\hat{H},\hat{H}^{*}\}$ were applied through three successive unitary transformations. At the output, the occurrence probabilities of the states $|1;0\rangle_{ab}$ and $|0;1\rangle_{ab}$ were measured from the photon number distribution obtained over a 1-second integration period. Using approximately 85,000 detection events, the correlation function with respect to the target state was calculated.
Fig. 3 shows a histogram of the obtained correlation function. To construct the histogram, the approximately 85,000 experimental data points were divided into subsets of 1,000 events each. The correlation function was calculated for each subset, and the frequency distribution of these values was plotted. From the histogram, the mean of the correlation function was 0.490, and the square root of the variance (standard error of the mean) was 0.017. These results indicate that the experimentally obtained correlation function significantly exceeds the limit of 0.25 achievable by the classical processor with one bit communication. In the figure the vertical blue line corresponds to this limit. Its value is derived in the next section.
Next, for the implementation of the one-qutrit processor, we attenuated a laser beam to the single-photon level and approximated a two-photon state by post-selecting events in which two photons were detected simultaneously (Fig. 2(c)). The total number of two-photon detection counts observed in this experiment was approximately 5,000 per measurement.
The input state for the qutrit experiment was $|2;0\rangle_{ab}$, corresponding to two photons in mode $a$. Four unitary transformations, randomly selected from the operations defined in Eqs.  and (1)–(4) and (7)–(10), were sequentially applied. The output probabilities of the states $|2;0\rangle_{ab}$ and $|0;2\rangle_{ab}$ were measured, and approximately 65,000 detection events were collected. Correlation functions were then calculated based on these data.
Fig. 4 shows a histogram of the correlation function for four module quantum processor with one qutrit communication. As in Fig. 3, the approximately 65,000 experimental data points were divided into subsets of 1,000 events each, and the correlation function was calculated for each subset. The histogram displays the frequency distribution of the resulting correlation values with respect to the target function (11). From the histogram, the mean of the correlation function was 0.525, and the square root of the variance was 0.014. These results indicate that the experimentally obtained correlation function significantly exceeds the classical limit which, as proven in the next section and shown in Fig. 4, is 0.375.
Figure 3: The histogram of the correlation function based on the data from 3-module single qubit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation which for the considered target function is equal to 0.25.
Figure 4: The histogram of the correlation function based on the data from 4-module single qutrit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation which for the considered target function is equal to 0.375.
Figure 5: The histogram of the correlation function based on the data from 3-module single qutrit processor experiment on silicon photonics. The blue line shows the classical limit for the correlation which for the considered target function is equal to 0.4375.
Finally, in post-processing, from the data of the qutrit experiment with four modules we extracted data which correspond to the first gate fixed as the identity. In this way we performed the analysis of the three-module processor with a qutrit communication. The target function in this case is given by the first four rows of (11). We built the histogram of the values or correlation function for 65 subsets of data with approximately 250 points in each subset. The average is 0.509 with the standard error of the average 0.068. This allows us to conclude that the observed value of the correlation significantly exceeds the limit for the correlation achievable by the classical processor with one-trit communication which, as proven in the next section and shown in Fig. 5, is equal to 0.4375.
In conclusion, in each case which we analyzed, we observe that the realistic experiment produces the correlation function close to 0.5, which is the maximum value of the correlation function with respect to the target (5) or (11) with half of the entries equal to zero. In Section IV we establish bounds on the achievable correlation function for optimized classical sequential processors which are shown in Figs. 3, 4, and 5 as the vertical blue lines. We conclude that the classical processor cannot output a function that approaches the target function for any deterministic strategy. In this sense, the quantum processor significantly outperforms the classical one in this task, showing a separation of expressivity between the classical and quantum processors.
IV Classical bounds on correlations
In this section, we establish bounds on the correlations with the target functions given in (5), (11) which are achievable by outputs of sequential classical processors with one-bit and one-trit communication, respectively, as shown in Figs. 3, 5, 4. The maximum correlation is uniquely defined by the minimum number of errors between the output of the classical processor and the target function. In the following parts, we establish these results as propositions.
IV.1 Bound for one-bit processor
Proposition 1.
The number of errors between the output of the three-stage processor with one-bit communication and the nonzero part of the target function (5) cannot be less than 8.
Proof.
In the proof, we explicitly use the properties of the target function (5). The rows of this matrix correspond to the processor action depending on the first-stage output. As the first module can have only two different symbols, the target reshuffled as a $4\times 16$ matrix needs to be rank 2. So, two lines need to be identical. Let these lines be 1st and 3rd (the same can be repeated for all other choices). We add the two rows, obtaining a binary vector of length 16.
We reshape this vector into a matrix $4\times 4$. It needs to be rank 2, as the rows describe the processor’s actions depending on the output from the second stage, which can output at most two symbols. However, we observe that each row is different and that the minimum number of different bits between each pair of rows (minimum Hamming distance) is 2. We need to correct two bits in two rows. In this way, we necessarily introduce 4 errors.
Finally, we have the remaining lines 2 and 4 of the target reshaped into the matrix $4\times 16$. When we reshape each of these lines into $4\times 4$ matrices, for the same reason as before, we need to correct 4 additional bits.
This justifies that the number of errors that we need to introduce to the target function (5) to make it compatible with the class of functions output by the discussed classical sequential processor is not less than 8.
Performing simulated annealing combinatorial optimization on Fixstars Amplify (details of this procedure are provided in the next section) we find a strategy for each module of the processor such that the final output approximates the non-zero part of the target with exactly 8 errors. The example is shown in Table 1.
$i$ $F_{i-1}=0$ $F_{i-1}=1$
$F_{1}$ (0,1,1,1) -
$F_{2}$ (0,1,1,0) (1,0,0,1)
$F_{3}$ (-1,1,-1,1) (1,-1,1,-1)
Table 1: Optimal strategies for one-bit processor with three nodes.
This concludes the proof. The minimum 8 errors correspond to the maximum value of the correlation between the output and the target equal to 0.25. ∎
Actually, because the size of the problem is relatively small, we could guess the solution with only 8 errors. However, we perform the simulation on the annealer to demonstrate that our numerical optimization tool works well.
IV.2 Bound for three-module trit processor
Proposition 2.
The number of errors between the output of the three-stage processor with one-trit communication and the nonzero part of the target function (11) (lines 1-4) cannot be less than 2.
Proof.
The target function is the first four rows of (11). As the rows of the matrix correspond to the functions in modules 2 and 3 given one of 3 symbols output from module 1, the rank of the matrix has to be equal to 3. Because the matrix is binary, the rank restriction requires that two of the rows be identical.
Without loss of generality, assume that the first and third rows are identical (the procedure can be repeated in all other cases). We complete the undefined elements of the first row by the defined elements of the third row. We observe that we have four blocks of size 4. Moreover, all are different, and the number of different bits between each pair of the four-blocks is equal to 2. We can reduce the number of these 4-size blocks to three by changing at least 2 bits. So, the smallest number of changes we would need is 2.
We can give an explicit example of a processor which approximate the target function with only two errors, as Table 2.
$i$ $F_{i-1}=0$ $F_{i-1}=1$ $F_{i-1}=2$
$F_{1}$ (0,1,2,1) - -
$F_{2}$ (2,1,2,1) (1,2,0,2) (2,1,1,2)
$F_{3}$ (-1,1,-1,1) (1,-1,-1,1) (-1,1,1,-1)
Table 2: Optimal strategies for one trit processor with three nodes.
This completes the proof of the proposition. The two errors correspond to the maximum value of the correlation function of 0.4375.
IV.3 Bound for four-module trit processor
Proposition 3.
The number of errors between the output of the four-stage processor with one-trit communication and the non-zero part of the target function (11) cannot be less than 16.
Proof.
The proof uses the properties of the target function (11) explicitly.
We reshape the target into the $4\times 64$ matrix. The rows correspond to the processor action depending on the first-stage output. As it can have only three different symbols, the target reshuffled as a $4\times 64$ matrix must be rank 3. So, two lines need to be identical. Let these lines be the first and second (the same can be repeated for all other choices). We add the two rows, obtaining a binary function of length 64.
We reshape this vector into a matrix $4\times 16$. It needs to be rank 3, as the rows describe the processor’s actions depending on the output from the second stage, which can contain at most 3 symbols. So, two rows of this matrix have to be the same. However, we observe that the minimum number of different bits between each pair or rows (minimum Hamming distance) is 8. Correcting some entries in one of the rows, we necessarily introduce 8 errors in the original target function.
We observe that the remaining lines of the $4\times 16$ matrix contain 4 different blocks of length 4. To reduce this number to 3 different blocks of length 4 we need to introduce 2 additional errors per line. This means that 4 additional errors are required to correct these parts.
Finally, we have the remaining lines 3 and 4 of the target reshaped into the $4\times 64$ matrix. When we reshape each of these lines into $4\times 16$ matrices, for the same reason as before, we expect that each of them is rank 3. So, the two lines need to be identical. We observe that when we add two complementary lines we get vectors of length 16 consisting of 4 blocks of size 4. To reduce this number to 3 we need to introduce 2 errors per each of the two rows.
This justifies that the number of errors that we need to introduce to the target function to make it compatible with the class of functions output by the discussed classical sequential processor is not smaller than 16.
Performing simulated annealing combinatorial optimization on Fixstars Amplify (details of this procedure are provided in the next section) we find an optimal strategy for each module of the processor such that the final output approximates the non-zero part of the target with exactly 16 errors. The example is shown in Table 3.
$i$ $F_{i-1}=0$ $F_{i-1}=1$ $F_{i-1}=2$
$F_{1}$ (0,2,1,0) - -
$F_{2}$ (1,2,0,1) (2,1,2,0) (2,0,1,2)
$F_{3}$ (1,0,1,2) (2,1,1,0) (1,2,0,1)
$F_{4}$ (1,-1,-1,1) (-1,1,1,-1) (-1,1,-1,1)
Table 3: Optimal strategies for one trit processor with four nodes.
This concludes the proof. The minimum number of errors 16 correspond to the maximum value of the correlation function equal to 0.375. ∎
In this section, we considered the problem of how well a classical sequential processor with a single bit or single trit communication can approximate a given target function of distributed input data. The target functions considered by now were specific ones generated by the outputs of quantum processors for the same data. By showing the minimum number of errors in the approximation, we show the gap in expressivity of the classical processor with respect to the quantum processor for specific tasks. The proofs of this section are based on explicit properties of the target functions.
In the next section, we consider a more general problem. Given an arbitrary binary target function, what is the best approximation by an output of a sequential processor? In general, this problem belongs to the high-complexity class. Solving it efficiently would imply an efficient algorithm for, for example, a low-rank binary matrix approximation, which is known to be NP-hard [27, 28, 29]. In the following section, we reduce the problem to an Ising problem, which can be approachable by nowadays and future Ising solvers based on simulated or quantum annealing, or other techniques.
V Optimal classical sequential processors
Consider an $N$ stage sequential processor with one-bit communication between modules and two-bit inputs in each module. The local inputs are not known to other modules. The performance of the processor can be optimized on the strategies of the modules.
We formulate two tasks that are addressed in the subsequent sections:
  1. 1.
    Decide if a given binary target function $T(\{X_{i}\})$ can be achieved as an output from the processor.
  2. 2.
    Determine the smallest Hamming distance between the output of the classical processor $O(\{X_{i}\})$ and the target function $T(\{X_{i}\})$.
V.1 Task 1: Existence criterion
For simplicity, assume that the number of processing modules is $N=3$. We can answer the first challenge by checking the number of different blocks of size 4, and 16 in the target function. There are 16 blocks of size 4. They are indexed by sequences of bits $(X_{1}X_{2}X_{3}X_{4})$ and defined as
$t^{[4]}(X_{1}...X_{4},X_{5}X_{6})=T(X_{1},...,X_{6}).$
There are 4 blocks of size 16. They are indexed by sequences of bits $(X_{1}X_{2})$ defined as
$t^{[16]}(X_{1}X_{2},X_{3}...X_{6})=T(X_{1},...,X_{6})$
Theorem 1. The processor defined above can output functions $O(X_{1},...,X_{6})$ with at most two different rows in the matrix $O^{[4]}(X_{1}X_{2},X_{3}...X_{6})$ and at most two different rows in the reshuffled matrix $O^{[16]}(X_{1}...X_{4},X_{5}X_{6})$.
Proof.
There are four possible outputs (for each $X_{1}X_{2}$, one) from the first processor but only two binary symbols 0 or $1$. For each symbol there is one and only one succession (the row of the matrix $O^{[4]}(X_{1}X_{2},X_{3}...X_{6})$), as the following processor cannot distinguish the same symbols output for different inputs of the first processor. So, there can be only two different rows in matrix $O^{[4]}$. Similarly, there are 16 outputs (for each sequence $(X_{1}...X_{4})$, one) from the second processor, but only two binary symbols. For each symbol, there is one succession. The row of the matrix $O^{[16]}(X_{1}...X_{4},X_{5}X_{6})$ for a chosen value of $(X_{1}...X_{4})$ is equal to one of these successions. If we recognize that the target function consists of only two distinct rows $O^{[16]}$ and only two distinct rows $O^{[4]}$ we can associate symbols 0 or $1$ as the output of the first processor in the positions corresponding to different distinct rows. Similarly, we associate 0 or $1$ as the outputs of the second processor in the positions corresponding to distinct rows of $O^{[16]}$. ∎
The argument can be easily extended to processors with more modules and different limits on communication between the modules.
Therefore, we can formulate the exisence criterion as the following theorem.
Theorem 2. [Existence criterion] For a binary function of several variables, one can check if it can be an output from a sequential processor with a single bit communication and a given distribution of variables between the modules by checking if each reshuffling of this function into a matrix with rows indexed by the local variables of earlier modules, and columns indexed by the remaining variables is row-rank at most 2.
For sequential processors with a single trit communication, the row-rank should be 3, etc.
V.2 Task 2: Optimal circuit as the Ising problem
V.2.1 Example
Figure 6: Two stage sequential processor with one bit communication and one bit input $X_{i}$ at each stage. The output function of inputs is denoted by $O(X_{1},X_{2})$.
For simplicity, let us consider an example – a two-module sequential processor with one bit communication and one bit input at each stage, as shown in Fig. 6. The first part, depending on the bit $X_{1}$ it obtains as the input, chooses which binary variable to pass to the next stage. Its strategy is given by a binary function $F_{1}$ which is defined by binary variables $a_{1}$ and $a_{2}$ returned for the respective input $X_{1}$. The second module strategy $F_{2}$ is a function of the bit passed from the previous part and one extra bit $X_{2}$. This function is given by a four-bit binary vector of binary variables $a_{3},...,a_{6}$, see Fig. 7. The goal is to choose the values of the variables $a_{1},...,a_{6}$ such that the output from the last part of the processor for given variables $X_{1}$ and $X_{2}$ was as close as possible to the given binary target function $T(X_{1},X_{2})$.
Figure 7: Example of binary variables which define each part’s strategy functions $F_{i}$ of their respective inputs. In this there are two parts in the processor. Each part processes the bit obtained from the previous part and one extra bit $X_{i}$.
For a two-stage processor, the problem is trivial because the target has only 4 values. We can associate the first two elements of the target function with $a_{3}$ and $a_{4}$, and the other two with $a_{5}$ and $a_{6}$ and put $a_{1}=0$ and $a_{2}=1$. However, we want to show that we can obtain this result as the optimization of an Ising-like problem. Then we generalize the entire formalism for any number of stages and more general setups in which the problem is highly nontrivial.
For any $X_{1}$ and $X_{2}$ the output of the processor $O(X_{1},X_{2})$ is
$O(X_{1},X_{2})=(1-a_{1+X_{1}})a_{3+X_{2}}+a_{1+X_{1}}a_{5+X_{2}}$ (12)
For binary functions, the minimum Hamming distance coincides with the maximum correlation function, which in this example is given as
$C$ $=$ $\sum_{\{X_{i}\}}T(X_{1},X_{2})O(X_{1},X_{2})$
$=$ $T(0,0)[(1-a_{1})a_{3}+a_{1}a_{5}]$
$+$ $T(0,1)[(1-a_{1})a_{4}+a_{1}a_{6}]$
$+$ $T(1,0)[(1-a_{2})a_{3}+a_{2}a_{5}]$
$+$ $T(1,1)[(1-a_{2})a_{4}+a_{2}a_{6}]$
The last formula is a quadratic form in binary variables, which can be written as an Ising problem with non-local two-body interactions. Its maximum provides the set of bits $a_{1},...,a_{6}$ that maximizes $C$.
V.2.2 Generalization
The construction of the example above can easily be generalized to arbitrary processor length, arbitrary communication, and arbitrary number of local input bits. Let us give here the explicit formula for a three-stage processor with a one-bit communication and a two-bit local input, as in the lower part of Fig. 3. The binary variables defining the strategies are shown in Fig. 8.
Figure 8: Example of binary variables which define each part’s strategy functions $F_{i}$ of their respective inputs. In this example there are three modules in the processor. Each part processes a bit obtained from the previous part and two extra bits.
The formula for the corelations between the target and the output reads
$C$ $=$ $\sum_{\{X_{i}\}}T(\{X_{i}\})$
$*$ $[(1-a_{1+2X_{1}+X_{2}})(1-a_{5+2X_{3}+X_{4}})a_{13+2X_{5}+X_{6}}$
$+$ $(1-a_{1+2X_{1}+X_{2}})a_{5+2X_{3}+X_{4}}a_{17+2X_{5}+X_{6}}$
$+$ $a_{1+2X_{1}+X_{2}}(1-a_{9+2X_{3}+X_{4}})a_{13+2X_{5}+X_{6}}$
$+$ $a_{1+2X_{1}+X_{2}}a_{9+2X_{3}+X_{4}}a_{17+2X_{5}+X_{6}}]$
In this case, the correlation function is a polynomial of degree three of binary variables which can be mapped in a standard way to the Ising model with up to three spins and possibly far-range interactions.
Further generalizations are needed to consider four-stage processors and the processor which allow for trit-based communication. In this case, variables $a_{i}$ are integer numbers from the set $\{0,1,2\}$. Because routinely we use solvers acting on binary variables, before formulating the formulas for correlations we first map each trit-variable to two one-bit variables. We do not show the final formulas that can be easily obtained with the same strategies as described above.
V.2.3 Solvers
In the previous sections, we show that finding the highest correlations between the output of the classical sequential processor with restricted communication and a given binary target function is equivalent to solving an Ising problem for a chain of spins with nonlocal and multi-spin interactions. This problem is, in general, NP-complete, and efficient solvers for every case do not exist. One uses approximate solvers applying specific heuristics.
In our research, we used a dedicated computer provided by Fixstars Amplify [30]. It finds solutions through the process of simulated annealing. We used the solver to find candidate solutions for the maximum correlation between the output of the classical three-module sequential processor with two-bit local variables and one-bit communication and the target given by the deterministic part of the output of the corresponding quantum processor with one-qubit communication. The problem involved 20 binary variables (spins) as shown in Fig. (8). We defined the Ising Hamiltonian corresponding to the correlation function (V.2.2). The solver was able to find the solution with 8 errors with respect to the target function. It is shown in Table 1. Proposition 1 states that this is the lower bound on the number of errors for this target. Therefore, we conclude that the solver found the optimal solution.
We repeated the numerical experiment on Fixstars Amplify solver successfully finding cases corresponding to the optimal strategies of the classical processors with a single trit communication, three- and four-modules and the target functions given by the deterministic parts of the corresponding quantum optical processors with qutrit communication as in the quantum optics experiment. The target functions are given (11) or the first four rows of it. The lower bounds on the number of errors are given in Propositions 2 and 3. As the optical processor could deterministically achieve only two solutions, $|20\rangle$ and $|02\rangle$, the number of bits in the output of the last stage was the same as in the bit processor, which is 4 for strategies associated with possibly three different outputs from the previous module. The total number of binary variables in the optimization of the three-module trit processor was 44. We found the strategies of each module that achieved the minimum number of errors which is shown in Table 2.
For finding the optimal four-module trit processor strategy with Fixstars Amplify, we needed (for our license restriction) to limit the number of binary variables to 48. We did it by guessing the strategies of the first and last modules. The solver found that the optimal strategies of the intermediate modules as shown in Table 3.
VI Extensions and Outlook
In this paper, we have studied theoretically and experimentally the separation between performance of quantum and classical processors of the same structure tackling the problem of approximation of certain functions of input variables when the variables are available only locally to different modules and the communication between the modules is restricted. In particular, binary functions can be achieved by the quantum sequential processors with qubit or qutrit communication, which can only be approximated by the output of the corresponding classical processors. We showed that existing silicon optics technology allows us to demonstrate this separation.
To prove the separation in the cases tailored to our experiment, we needed to develop a general theoretical framework which allows us to put strict boundaries on the performance of classical processors. We did it two ways. From the properties of specific target binary functions related to the output of our experiment, and from the reduction of the optimization in the set of binary function which some properties to the Ising problem for arbitrary binary targets.
If the target function is an arbitrary binary function, the Ising optimization, or alternatively the optimization of the corresponding correlation function gives the solution to the problem of the best approximation of one arbitrary binary function by another which satisfies some restrictions. In our case, the restriction is to be the output of a sequential processor with communication between the modules restricted to a link with given dimensionality. The restriction has a clear interpretation when we reshape the function in the form of a matrix, the rows of which correspond to the processor response to given outputs of one module. As the number of symbols in the output is restricted, the number of different rows is restricted. The number of different rows in a binary matrix corresponds to the rank. So, the problem of finding the optimal output of a two-module processor with limited communication to an arbitrary binary function directly corresponds to the problem of approximation of an arbitrary binary matrix by a binary matrix of specific rank. The optimality is defined here by the Hamming distance.
Unlike in low rank approximation of real or complex matrices, which can be done by reducing rank in the singular value decomposition, the low rank approximation of binary matrices is known as the hard problem. A side result of our work is the reduction of this problem to the generalized Ising problem, which is also known to be hard in general, but dedicated approximate solvers exist and are still improving.
Another restriction on approximation of the general binary target function of a number of variables can be the output of a multi-module processor. In this case, we approximate a general tensor (indexed by the variables) in terms of a concatenation of tensors of specific ranks. Finding an optimal approximation in terms of the Hamming distance is again a difficult task, which we explicitly reduce to a generalized Ising problem and can attack with existing Ising solvers.
The third related application of the techniques we developed in this paper is the binary matrix completion. We could ask, what is a good fixed rank binary matrix completion of a binary matrix with restricted number of known entries. Again we tacle this problem by creating the binary matrix with highest possible correlation function with the existing part of the reference matrix and with required values of the rank or ranks in different matrix reshapings.
Finally, the direct application of our techniques is the logic and sequential logic design. In fact, given an arbitrary binary target function of a number of variables can be interpreted as the truth table for some logical formulas of the corresponding Boolean variables. The optimal construction of the logical circuits or circuits satisfying specific conditions such as being distributed or sequential is a highly desirable and unsolved problem. Although approximations exist and are widely used, new techniques, such as the one proposed in this paper, can contribute to the design improvement.
Acknowledgements.
We would like to thank Marek Żukowski for valuable discussions. This work was supported by JST Moonshot R&DR\&D, Grant No. JPMJMS226C and Grant No. JPMJMS2061, JST ASPIRE, Grant No. JPMJAP2427, JST COI-NEXT Grant No. JPMJPF2221, JST ERATO Grant No. JPMJER2402, and JSPS KAKENHI Grant Number JP24K00559.